Which values of k and t can make y'=a*(x^k)+b*(y^t) a homogeneous equation with the help of y=z^m
1
Expert's answer
2020-12-15T07:09:20-0500

If "y=x^m" , then:

"mx^{m-1}=ax^k+bx^{mt}"

Condition for a homogeneous equation:

"m-1=k=mt"

So:

"k=m-1, t=\frac{m-1}{m}"